How many moles of hydrogen are in 1.4 moles of quinine?
1
Expert's answer
2016-12-09T03:08:08-0500
Quinine - C20H24N2O2
n(H) = 24* n(C20H24N2O2) = 24*1.4 = 33.6 (mol)
Answer
n(H) = 33.6 (mol